OpenStudy (anonymous):

Solve 4^(x + 7) = 6^(x – 1)

OpenStudy (amistre64):

might wanna log it out

OpenStudy (amistre64):

\[a^m=b^n\] \[ln(a^m)=ln(b^n)\] \[mln(a)=nln(b)\] \[mln(a)-nln(b)=0\] should be easeir to collect like terms that way